OpenStudy (amistre64):

In a vote, 1 = FOR, 2 = AGAINST. Estimate the probability that it would the vote passes. 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1; (12) votes 2, 2, 2, 2, 2, 2, 2, 2; (8) votes A) 0.6 B) 0.5 C) 0.65 D) 0.4

OpenStudy (anonymous):

A!

OpenStudy (anonymous):

60% of 20 is 12. :)

OpenStudy (amistre64):

yeah, i see that now that I formated it all like that :) is there a way to do this with nPr or nCr ?

OpenStudy (anonymous):

???? hmm

OpenStudy (anonymous):

The way I did it was % times of = is % times 20 = 12 12/20 = .6 = 60% :)

OpenStudy (anonymous):

I take it back, what's the threshold for the vote to pass. Unanimous,majority?

OpenStudy (amistre64):

doesnt say; this is what they gave me in the study guide... A polling firm, hired to estimate the likelihood of the passage of an up-coming referendum, obtained the set of survey responses to make its estimate. The encoding system for the data is: 1 = FOR, 2 = AGAINST. If the referendum were held today, estimate the probability that it would pass. 1, 2, 2, 1, 1, 2, 1, 2, 2, 1, 1, 1, 2, 1, 2, 1, 1, 1, 2, 1

OpenStudy (anonymous):

A!

OpenStudy (amistre64):

and for some reason i was stuck on the track that i needed to figure permutations and combonations..

OpenStudy (anonymous):

NOPE! % times of = is :)

OpenStudy (amistre64):

\[\large \frac{20!}{8!12!}\]wasnt cutting the mustard

OpenStudy (anonymous):

the vote would pass as long as yay is more than nay

OpenStudy (anonymous):

20 votes all together 12 is the most for pass But what percent of 20 is 12? So I did % times of = is x * 20 = 12 Divided by 20 on both sides! .6 = 60 %